# Lennart Berg and Torben Wiedemann

### v1

@@ 3 Root graded groups
Throughout this section, we denote by $\Phi$ an arbitrary root system with root base $\Delta$. From after 3.4 on, we denote by $(G,(U_{\alpha})_{\alpha\in\Phi})$ a $\Phi$-graded group.
In this section, we introduce root graded groups in the sense of [Wie23].
@@
Let $G$ be a group and let $(U_{\alpha})_{\alpha\in\Phi}$ be a family of non-trivial subgroups. We say that $G$ has $\Phi$-commutator relations with root groups $(U_{\alpha})_{\alpha\in\Phi}$ if it satisfies the following conditions:
  1. (i)
    $[U_{\alpha},U_{\beta}]\subseteq U_{\mathord{]\alpha,\beta[}}$ for all non-proportional $\alpha,\beta\in\Phi$.
  2. (ii)
    Let $\alpha\in\Phi$ and $\beta\in\mathbb{R}_{>0}\alpha\cap\Phi$. Then $[U_{\alpha},U_{\beta}]\subseteq U_{\alpha+\beta}$ if $\alpha+\beta$ is a root and $[U_{\alpha},U_{\beta}]=\{1_{G}\}$ otherwise. Further, $U_{\beta}\subseteq U_{\alpha}$ if $\beta\in\mathbb{Z}_{>0}\alpha\cap\Phi$.
If, moreover, we have $[U_{\alpha},U_{\beta}]\subseteq U_{\mathord{]\alpha,\beta[}_{\text{cry}}}$ for all non-proportional $\alpha,\beta\in\Phi$, we say that $G$ has crystallographic $\Phi$-commutator relations with root groups $(U_{\alpha})_{\alpha\in\Phi}$.
@@
Axiom 3.3 (ii) differs from the axiom (RGG-Div) used in [Wie23, 2.1.4, 2.5.2]. The reason for this is, as explained in [Wie23, 2.1.9], that (RGG-Div) is specifically tailored for root systems which are either reduced or crystallographic. The root system $GH_{3}$, which we will encounter in 4.5, satisfies neither assumption, so an adjustment of this axiom is necessary. If $\Phi$ is reduced, then a $\Phi$-grading in the sense of [Wie23, 2.5.2] is a $\Phi$-grading in the sense of 3.4 if and only if all root groups are abelian. If $\Phi=H_{3}$, we will see in 5.3 (i) that the latter assumption is automatically satisfied without requiring Axiom 3.3 (ii), so the two notions of $H_{3}$-graded groups coincide.
@@
The following assertion is a consequence of Axiom 3.4 (iv).
Proposition 3.10 ([Wie23, 2.5.9, 2.1.19]).
Assume that $\Phi$ is reduced, let $\beta,\gamma\in\Phi$ be non-proportional and denote by $(\alpha_{1},\ldots,\alpha_{k})$ the interval ordering of $\mathord{]\beta,\gamma[}$ starting from $\beta$ (as in 2.7). Then the product map
@@
Using 3.10, we can introduce the following notation, which will be used frequently.
Notation 3.11.
Assume that $\Phi$ is reduced, let $\beta,\gamma\in\Phi$ be non-proportional and denote by $(\alpha_{1},\ldots,\alpha_{k})$ the interval ordering of $\mathord{]\beta,\gamma[}$ starting from $\beta$ (as in 2.7). Let $x_{\beta}\in U_{\beta}$ and $x_{\gamma}\in U_{\gamma}$. By 3.10, there exist unique elements $x_{i}\in U_{\alpha_{i}}$ for all $i\in\{1,\ldots,k\}$ such that $[x_{\beta},x_{\gamma}]=x_{1}\cdots x_{k}$. We put $[x_{\beta},x_{\gamma}]_{\alpha_{i}}\coloneqq x_{i}$ for all $i\in\{1,\ldots,n\}$. When we write $[x_{\beta},x_{\gamma}]_{\alpha_{i}}^{g}$ for some $g\in G$ or $[x_{\beta},x_{\gamma}]_{\alpha_{i}}^{-1}$, we always mean $([x_{\beta},x_{\gamma}]_{\alpha_{i}})^{g}$ or $([x_{\beta},x_{\gamma}]_{\alpha_{i}})^{-1}$, respectively.
@@
Lemma 3.13 ([Wie23, 2.1.18, 2.2.9]).
Assume that $\Phi$ is reduced, let $\alpha,\gamma\in\Phi$ be non-proportional, let $\beta\in\mathord{]\alpha,\gamma[}$ and let $x_{\alpha}\in U_{\alpha}$, $x_{\gamma}\in U_{\gamma}$. Then for any root $\delta$ and any $\delta$-Weyl element $w_{\delta}$, we have
@@
Remark 3.14.
The assertion that $[x_{\alpha},x_{\gamma}]_{\beta}^{-1}=[x_{\gamma},x_{\alpha}]_{\beta}$ is stated in [Wie23, 2.1.18] only for the case that $\mathord{]\alpha,\gamma[}$ has at most two elements, but the same proof remains valid in general. This fact is related to 3.12.
The well-known braid relations in Coxeter groups also hold for Weyl elements in root graded groups. This result goes back to [TW02, (6.9)], where it is stated and proven in a less general context. However, essentially the same proof remains valid in our setting, as is verified in [Wie23, 2.2.34].
@@
Assume that $\Phi$ is reduced, let $\alpha,\beta\in\Delta$ and denote by $m$ the order of $\sigma_{\alpha}\sigma_{\beta}$ in the Weyl group. Then for all Weyl elements $w_{\alpha}\in M_{\alpha}$, $w_{\beta}\in M_{\beta}$, we have
@@ 4 Foldings
Let $n\geq 4$. In this section, we use the standard representations of the root systems $A_{n-1}$ and $D_{n}$ from [Hum78, 2.10]:
@@
Here $(e_{1},\ldots,e_{n})$ denotes the standard basis of $\mathbb{R}^{n}$ (with the standard inner product). Further, we fix a commutative ring $\mathcal{R}$.
@@
Given a root system $\Phi$ and an automorphism $\rho$ of its Dynkin diagram, one can construct a new root system $\Phi^{\prime}$ by projecting $\Phi$ orthogonally to the fixed space of an isometry $\tau$ which is related to $\rho$ (with $\tau=\rho$ if $\Phi$ is reduced, which is the only case we will be interested in). In particular, we have a map $\Phi\rightarrow\Phi^{\prime}$ which we call a folding, and we will also refer to $\Phi^{\prime}$ as a folding of $\Phi$. Further, any Chevalley group of type $\Phi$ gives rise to a so-called twisted Chevalley group of type $\Phi^{\prime}$. This procedure is described in [Car72, Chapter 13]. In a similar way, every $\Phi$-grading $(U_{\alpha})_{\alpha\in\Phi}$ of a group $G$ gives rise to a $\Phi^{\prime}$-grading $(U_{\alpha^{\prime}}^{\prime})_{\alpha^{\prime}\in\Phi^{\prime}}$ of (the same group) $G$.
It is well-known that there exists a more general class of foldings which do not arise from diagram automorphisms. See [Mü93] for a description of this phenomenon. The goal of this section is to construct an $H_{3}$-graded group $G$ from a Chevalley group of type $D_{6}$ by way of the folding $D_{6}\rightarrow H_{3}$. Further, we will compute the commutator relations in $G$ (Figure 4). The main result of this paper is that every $H_{3}$-graded group has a coordinatisation by $\mathcal{R}^{\prime}\times\mathcal{R}^{\prime}$ for a commutative ring $\mathcal{R}^{\prime}$ such that the same commutator relations as in $G$ are satisfied (Theorem 8.31) and that every $H_{3}$-graded group arises from a $D_{6}$-graded group by the folding procedure described in this section (Theorem 8.40).
To illustrate the general folding procedure, we begin with the easier example of the folding $A_{4}\rightarrow H_{2}$.
Notation 4.2.
In order to emphasise that we consider specific examples of root graded groups (and not arbitrary root graded groups) in this section, we place a tilde on the root groups. Further, for the purpose of distinguishing different gradings of the same surrounding group $G$, we will use superscripts to indicate the root system. For example, we will write $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$ for an $A_{4}$-grading.
@@
For $1\leq i\neq j\leq 5$ denote by $e_{ij}$ the $(5\times 5)$-matrix with $1_{\mathcal{R}}$ at position $(i,j)$ and $0_{\mathcal{R}}$ at every other position. For all roots $\alpha=e_{i}-e_{j}\in A_{4}$, we define a root group
$\tilde{U}_{\alpha}^{A}\coloneqq\tilde{U}_{ij}^{A}\coloneqq\{\mathord{\mathord{\mathrm{id}}}+ae_{ij}\mid a\in\mathcal{R}\}\leq\operatorname{SL}_{5}(\mathcal{R}).$
We denote by $\operatorname{E}_{5}(\mathcal{R})$ the group generated by these root groups, which is a Chevalley group of type $A_{4}$. Then $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$ is an $A_{4}$-grading of $\operatorname{E}_{5}(\mathcal{R})$.
@@
By a straightforward computation, which we leave out because this example serves only motivational purposes, $(\tilde{U}_{\chi}^{H})_{\chi\in H_{2}}$ is an $H_{2}$-grading of $\operatorname{E}_{5}(\mathcal{R})$. We call $(\tilde{U}_{\chi}^{H})_{\chi\in H_{2}}$ a folding of $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$. Note that the surrounding group $\operatorname{E}_{5}(\mathcal{R})$ is the same for both gradings: only the family of root groups changes.
Observe that the root group of the simple root $\chi_{1}$ is obtained by “merging” the root groups of the simple roots $\delta_{1}\coloneqq e_{1}-e_{2}$ and $\delta_{3}\coloneqq e_{3}-e_{4}$ while $\tilde{U}_{\chi_{5}}^{H}$ is obtained from the simple roots $\delta_{2}\coloneqq e_{2}-e_{3}$ and $\delta_{4}\coloneqq e_{4}-e_{5}$. Hence our folding corresponds to the partition $\{\{\delta_{1},\delta_{3}\},\{\delta_{2},\delta_{4}\}\}$ of the root base $\{\delta_{1},\ldots,\delta_{4}\}$ of $A_{4}$, which we encode by the diagram in Figure 1(a). Note further that for each $i\in\{1,\ldots,10\}$, the two root groups in the definition of $\tilde{U}_{i}^{H}$ commute, so that $\tilde{U}_{i}^{H}$ is isomorphic to $\mathcal{R}\times\mathcal{R}$.
$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{4}$
(a) The folding $A_{4}\rightarrow H_{2}$.
$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{6}$$\delta_{4}$$\delta_{5}$
(b) The folding $D_{6}\rightarrow H_{3}$.
$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{4}$$\delta_{5}$$\delta_{6}$$\delta_{7}$$\delta_{8}$
(c) The folding $E_{8}\rightarrow H_{4}$.
Figure 1: Three foldings.

### v2

@@ 3 Root graded groups
Throughout this section, we denote by $\Phi$ an arbitrary reduced root system with root base $\Delta$. From after 3.5 on, we denote by $(G,(U_{\alpha})_{\alpha\in\Phi})$ a $\Phi$-graded group.
In this section, we introduce $\Phi$-graded groups in the sense of [Wie24, 2.5.2]. They generalise Tits’ notion of RGD-systems (see 7.10 and [Tit92, p. 258]).
@@
Let $G$ be a group and let $(U_{\alpha})_{\alpha\in\Phi}$ be a family of non-trivial subgroups. We say that $G$ has $\Phi$-commutator relations with root groups $(U_{\alpha})_{\alpha\in\Phi}$ if $[U_{\alpha},U_{\beta}]\subseteq U_{\mathord{]\alpha,\beta[}}$ for all non-proportional $\alpha,\beta\in\Phi$. If, moreover, we have $[U_{\alpha},U_{\beta}]\subseteq U_{\mathord{]\alpha,\beta[}_{\text{cry}}}$ for all non-proportional $\alpha,\beta\in\Phi$, we say that $G$ has crystallographic $\Phi$-commutator relations with root groups $(U_{\alpha})_{\alpha\in\Phi}$.
@@
Definition 3.5.
Let $(G,(U_{\alpha})_{\alpha\in\Phi})$ a $\Phi$-graded group and let $\alpha\in\Phi$. A root isomorphism (for $\alpha$) is an isomorphism $\theta_{\alpha}\colon M\rightarrow U_{\alpha}$ for a group $M$. Given such an isomorphism, we also say that $U_{\alpha}$ is parametrised by $M$.
@@
Note that the root groups in a $\Phi$-graded group are not assumed to be abelian. However, we will see in 5.3 (i) that $U_{\alpha}$ is necessarily abelian if $\alpha$ is contained in an $A_{2}$-subsystem. In particular, root gradings of types $H_{3}$ and $H_{4}$ have abelian root groups.
@@
The following two assertions are consequences of Axiom 3.4 (iv). The second, more general assertion will only be needed in 7.5 and 7.8.
Proposition 3.11 ([Wie24, 2.5.9, 2.1.19]).
Let $\beta,\gamma\in\Phi$ be non-proportional and denote by $(\alpha_{1},\ldots,\alpha_{k})$ the interval ordering of $\mathord{]\beta,\gamma[}$ starting from $\beta$ (as in 2.12). Then the product map
@@
Proposition 3.12.
Let $\Pi$ be a positive system in $\Phi$ and put $k\coloneqq\lvert\Pi\rvert$. Then there exists an ordering $\alpha_{1},\ldots,\alpha_{k}$ of the roots in $\Pi$ such that for every group $G^{\prime}$ which has $\Phi$-commutator relations with root groups $(U_{\alpha}^{\prime})_{\alpha\in\Phi}$, the following properties hold:
  1. (a)
    The product map
    $U_{\alpha_{1}}^{\prime}\times\cdots\times U_{\alpha_{k}}^{\prime}\rightarrow U_{\Pi}^{\prime},(x_{1},\ldots,x_{k})\mapsto x_{1}\cdots x_{k}$
    is surjective.
  2. (b)
    If $(U_{\alpha}^{\prime})_{\alpha\in\Phi}$ satisfies Axiom 3.4 (iv), then the product map above is bijective.
Proof.
By [Wie24, 2.3.23], there exists an ordering $\alpha_{1},\ldots,\alpha_{k}$ which is extremal in the sense of [Wie24, 2.3.16]. For any such ordering, the product map is surjective by [Wie24, 2.3.21, 2.4.14]. By [Wie24, 2.4.17], the product map is even bijective for extremal orderings if Axiom 3.4 (iv) is satisfied. ∎
Using 3.11, we can introduce the following notation, which will be used frequently.
Notation 3.13.
Let $\beta,\gamma\in\Phi$ be non-proportional and denote by $(\alpha_{1},\ldots,\alpha_{k})$ the interval ordering of $\mathord{]\beta,\gamma[}$ starting from $\beta$ (as in 2.12). Let $x_{\beta}\in U_{\beta}$ and $x_{\gamma}\in U_{\gamma}$. By 3.11, there exist unique elements $x_{i}\in U_{\alpha_{i}}$ for all $i\in\{1,\ldots,k\}$ such that $[x_{\beta},x_{\gamma}]=x_{1}\cdots x_{k}$. We put $[x_{\beta},x_{\gamma}]_{\alpha_{i}}\coloneqq x_{i}$ for all $i\in\{1,\ldots,n\}$. When we write $[x_{\beta},x_{\gamma}]_{\alpha_{i}}^{g}$ for some $g\in G$ or $[x_{\beta},x_{\gamma}]_{\alpha_{i}}^{-1}$, we always mean $([x_{\beta},x_{\gamma}]_{\alpha_{i}})^{g}$ or $([x_{\beta},x_{\gamma}]_{\alpha_{i}})^{-1}$, respectively.
@@
Lemma 3.15 ([Wie24, 2.1.18, 2.2.9]).
Let $\alpha,\gamma\in\Phi$ be non-proportional, let $\beta\in\mathord{]\alpha,\gamma[}$ and let $x_{\alpha}\in U_{\alpha}$, $x_{\gamma}\in U_{\gamma}$. Then for any root $\delta$ and any $\delta$-Weyl element $w_{\delta}$, we have
@@
Note 3.16.
The assertion that $[x_{\alpha},x_{\gamma}]_{\beta}^{-1}=[x_{\gamma},x_{\alpha}]_{\beta}$ is stated in [Wie24, 2.1.18] only for the case that $\mathord{]\alpha,\gamma[}$ has at most two elements, but the same proof remains valid in general. This fact is related to 3.14.
The following notation is a variation of 3.13.
Definition 3.17.
Assume that $(G,(U_{\alpha})_{\alpha\in\Phi})$ is equipped with a family of root isomorphisms $(\theta_{\alpha}^{X}\colon M\rightarrow U_{\alpha}^{X})_{\alpha\in X}$ for a group $M$. For all non-proportional roots $\alpha,\gamma\in\Phi$ and all $\beta\in\mathord{]\alpha,\gamma[}$, we define a map $\psi_{\alpha,\gamma}^{\beta}\colon M\times M\rightarrow M$ by the property that
for all $x,y\in M$. It is called the commutation map of $G$ at position $(\alpha,\gamma,\beta)$ (with respect to $(\theta_{\delta})_{\delta\in\Phi}$). If $\mathord{]\alpha,\gamma[}$ has only one element $\beta$, we will write $\psi_{\alpha,\gamma}$ in place of $\psi_{\alpha,\gamma}^{\beta}$.
The well-known braid relations in Coxeter groups also hold for Weyl elements in root graded groups. This result goes back to [TW02, (6.9)], where it is stated and proven in a less general context. However, essentially the same proof remains valid in our setting, as is verified in [Wie24, 2.2.34].
@@
Let $\alpha,\beta\in\Delta$ and denote by $m$ the order of $\sigma_{\alpha}\sigma_{\beta}$ in the Weyl group. Then for all Weyl elements $w_{\alpha}\in M_{\alpha}$, $w_{\beta}\in M_{\beta}$, we have
@@ 4 Foldings
In this section, we consider the root systems $A_{4}$, $D_{6}$ and $E_{8}$ with the labelling of simple roots given by Figure 2. Note that our ordering of simple roots in $E_{8}$ is non-standard: It is chosen so that we can identify the simple roots $\delta_{1},\ldots,\delta_{6}$ in $D_{6}$ with the simple roots $\delta_{1},\ldots,\delta_{6}$ in $E_{8}$. We will also use the standard representations
@@
where $n\geq 4$ and $(e_{1},\ldots,e_{n})$ denotes the standard basis of $\mathbb{R}^{n}$ (with the standard inner product). Further, we fix a root base $\Delta_{H}=(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$ of $H_{4}$ in standard order and we make the identifications $H_{2}=\Psi_{\rho_{2},\rho_{3}}$ and $H_{3}=\Psi_{\rho_{1},\rho_{2},\rho_{3}}$. We will also consider $H_{2}$, $H_{3}$ and $H_{4}$ as subsets of $GH_{2}$, $GH_{3}$ and $GH_{4}$.
Certain arguments and remarks will apply to all three root systems $A_{4}$, $D_{6}$ and $E_{8}$ simultaneously. In such situations, we simply write $X$ to denote an arbitrary root system $X\in\{A_{4},D_{6},E_{8}\}$. We also put $\ell\coloneqq\operatorname{rank}(X)/2$ (because we will construct a folding $X\rightarrow H_{\ell}$) and denote by $\Delta_{X}$ the ordered root base of $X$ given by Figure 2.
$\delta_{2}$$\delta_{3}$$\delta_{4}$$\delta_{5}$
(a) The folding $A_{4}\rightarrow H_{2}$.
$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{6}$$\delta_{4}$$\delta_{5}$
(b) The folding $D_{6}\rightarrow H_{3}$.
$\delta_{7}$$\delta_{1}$$\delta_{2}$$\delta_{3}$$\delta_{4}$$\delta_{6}$$\delta_{8}$$\delta_{5}$
(c) The folding $E_{8}\rightarrow H_{4}$.
Figure 2: Three foldings.
@@
Given a root system $\Phi$ and an automorphism $\rho$ of its Dynkin diagram, one can construct a new root system $\Phi^{\prime}$ by projecting $\Phi$ orthogonally to the fixed space of an isometry $\tau$ which is related to $\rho$ (with $\tau=\rho$ if $\Phi$ is simply-laced, which is the only case we will be interested in). In particular, we have a map $\Phi\rightarrow\Phi^{\prime}$ which we call a folding, and we will also refer to $\Phi^{\prime}$ as a folding of $\Phi$. Further, any Chevalley group of type $\Phi$ gives rise to a so-called twisted Chevalley group of type $\Phi^{\prime}$. This procedure is described in [Car72, Chapter 13]. In a similar way, every $\Phi$-grading $(U_{\alpha})_{\alpha\in\Phi}$ of a group $G$ gives rise to a $\Phi^{\prime}$-grading $(U_{\alpha^{\prime}}^{\prime})_{\alpha^{\prime}\in\Phi^{\prime}}$ of (the same group) $G$.
A more general class of foldings, which do not arise from diagram automorphisms, is described in [Mü93]. In this section, we illustrate how, in the notation of 4.1, the root system $H_{\ell}$ can be realised in this way as a folding of $X\in\{A_{4},D_{6},E_{8}\}$ (see 4.8). We will also describe a procedure to construct $H_{\ell}$-graded groups from $X$-graded groups in a way which generalises the aforementioned construction of twisted Chevalley groups (see 4.19 and 4.21). This general procedure can be applied to the example of Chevalley groups of type $X$, and we will compute the commutator relations in the foldings of these specific examples (see 4.29) as well as some more structure constants (see 4.27).
The main result of this section is 4.32, which shows that the folding procedure actually produces an $H_{\ell}$-grading in the sense of 3.4. Some technicalities of the proof are only dealt with in the appendix, which is independent of the following sections of this paper. However, even without the appendix, the arguments (and computations) in this section are sufficient to prove that the foldings of the aforementioned Chevalley groups are $H_{\ell}$-gradings.
Since the root groups of a Chevalley group (of type $X$) are parametrised by a commutative ring $\mathcal{R}$ (in the sense of 3.5), the root groups of the corresponding $H_{\ell}$-grading are parametrised by $\mathcal{R}\times\mathcal{R}$ in a natural way. Hence we say that the folding of a Chevalley group of type $X$ is coordinatised by $\mathcal{R}\times\mathcal{R}$, a notion that we will make precise in 6.28. The main result of this paper is that if $\ell\geq 3$, then every $H_{\ell}$-graded group has a coordinatisation by $\mathcal{R}\times\mathcal{R}$ for a commutative ring $\mathcal{R}$ (Theorem 9.33). This means, essentially, that the commutator relations in an arbitrary $H_{\ell}$ grading have the same form as in 4.29. It is a consequence of this result that for $\ell\geq 3$, every $H_{\ell}$-graded group arises from an $X$-graded group (but not necessarily from a Chevalley group, see 7.11) by the folding procedure described in this section (Theorem 10.11). Further, we point out that by our construction in this section, there exists an $H_{\ell}$-graded group coordinatised by $\mathcal{R}\times\mathcal{R}$ for every commutative ring $\mathcal{R}$ (and even for every ring $\mathcal{R}$ if $\ell=2$).
We will obtain the root system $GH_{\ell}$ as a natural intermediate step in the folding $X\rightarrow H_{\ell}$. Essentially, each root in $H_{\ell}$ is associated to two roots in $GH_{\ell}$, with short roots in $GH_{\ell}$ corresponding to the left factor of $\mathcal{R}\times\mathcal{R}$ in our coordinatisation and long roots corresponding to the right factor. Hence the root system $GH_{\ell}$ allows us to see the structure of an $H_{\ell}$-graded group $G$ more clearly if $G$ is obtained via folding (which, by the previous paragraph, is always the case for $\ell\geq 3$). See 4.16 and A.11 for more details.
We point out that our focus in this section lies on the cases $X=D_{6}$ and $X=E_{8}$. The case $X=A_{4}$ is included mainly because this higher generality does not introduce any additional difficulties.
Note 4.2.
While Chevalley groups over fields are well-known objects (see, for instance, [Car72, Ste16]), the notion of Chevalley groups over commutative rings is much less standard. For an introduction, see [VP96].
To be more precise, the objects that we call “Chevalley groups” are called “elementary Chevalley groups” in [VP96]. Chevalley groups of type $\Phi$ have a natural $\Phi$-grading $(U_{\alpha})_{\alpha\in\Phi}$. Specifically, (elementary) Chevalley groups are generated by their root groups by definition, they have $\Phi$-commutator relations by the Chevalley commutator formula and Weyl elements exist by [VP96, (13.1)]. While we are not aware of a reference which explicitly states Axiom 3.4 (iv) for Chevalley groups over rings, the proof in [Ste16, Corollary 3 of Lemma 17] for Chevalley groups over fields remains valid in this generality.
We also point out that we provide explicit matrix representation of (certain) Chevalley groups of types $A_{4}$ and $D_{6}$ in 4.3 and 4.22. Hence our treatment of these cases is self-contained.
As an introductory example, we begin with the construction of an $H_{2}$-graded group from a Chevalley group of type $A_{4}$.
@@
Let $\mathcal{R}$ be a commutative ring. For all $1\leq i\neq j\leq 5$, we denote by $e_{ij}$ the $(5\times 5)$-matrix with $1_{\mathcal{R}}$ at position $(i,j)$ and $0_{\mathcal{R}}$ at every other position, and we define a root group
$\tilde{U}_{\alpha}^{A}\coloneqq\tilde{U}_{ij}^{A}\coloneqq\{\mathord{\operatorname{id}}+ae_{ij}\mid a\in\mathcal{R}\}\leq\operatorname{SL}_{5}(\mathcal{R}).$
We denote by $\operatorname{Ch}(A_{4},\mathcal{R})$ the group generated by these root groups, which is a (simply connected) Chevalley group of type $A_{4}$. Then $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$ is an $A_{4}$-grading of $\operatorname{Ch}(A_{4},\mathcal{R})$.
@@
By a straightforward computation, which we leave out because this example serves only motivational purposes, $(\tilde{U}_{\chi}^{H})_{\chi\in H_{2}}$ is an $H_{2}$-grading of $\operatorname{E}_{5}(\mathcal{R})$. We call $(\tilde{U}_{\chi}^{H})_{\chi\in H_{2}}$ a folding of $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$. Note that the surrounding group $\operatorname{E}_{5}(\mathcal{R})$ is the same for both gradings: only the family of root groups changes. Note further that for each $i\in\{1,\ldots,10\}$, the two root groups in the definition of $\tilde{U}_{i}^{H}$ commute, so that $\tilde{U}_{i}^{H}$ is isomorphic to $\mathcal{R}\times\mathcal{R}$
To pave the way for a generalisation, we rephrase our construction in a more abstract way. Denote by
$\delta_{2}\coloneqq e_{1}-e_{2},\quad\delta_{3}\coloneqq e_{2}-e_{3},\quad\delta_{4}-e_{3}-e_{4},\quad\delta_{5}\coloneqq e_{4}-e_{5}$
a root base of $A_{4}$ and define a linear map
$\bar{\pi}\colon\langle A_{4}\rangle_{\mathbb{R}}\rightarrow\langle H_{2}\rangle,\delta_{2},\delta_{4}\mapsto\chi_{1},\delta_{3},\delta_{6}\mapsto\chi_{5}.$
Then a straightforward computation (which, again, we leave out) shows that $\bar{\pi}$ induces a surjection $\pi\colon A_{4}\rightarrow H_{2}$ with fibers of cardinality $2$, and we have
$\tilde{U}_{\beta}^{H}=\langle\tilde{U}_{\alpha}^{A}\mid\alpha\in\pi^{-1}(\beta)\rangle$
for all $\alpha\in H_{2}$. We encode this folding by the partition $\{\{\delta_{2},\delta_{4}\},\{\delta_{3},\delta_{5}\}\}$ of the root base $\{\delta_{2},\delta_{3},\delta_{4},\delta_{5}\}$. We also say that $(\tilde{U}_{\chi}^{H})_{\chi\in H_{2}}$ is the folding of $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$ described by the diagram in Figure 2(a). .
Notation 4.4.
The tilde on the root groups in 4.3 represents the fact that $(\tilde{U}_{\alpha}^{A})_{\alpha\in A_{4}}$ is a specific example of an $A_{4}$-grading and not an arbitrary $A_{4}$-grading. We will use a similar notation in the later Examples 4.22 and 4.23.
@@
Everything in 4.3 remains correct if the ring $\mathcal{R}$ is not assumed to be commutative, except that the resulting group is then no longer called a Chevalley group, and that there is no longer a standard notion of $\operatorname{SL}_{5}(\mathcal{R})$.
Remark 4.6.